# VC Search: Bridging the Gap Between Well-Defined and Ill-Defined Problems in Mathematical Reasoning

### v2

@@ Abstract
Shi-Yu Tian12†, Zhi Zhou1†, Kun-Yang Yu1,2, Ming Yang1,2, Lin-Han Jia1, Lan-Zhe Guo13‡, Yu-Feng Li12‡ 1National Key Laboratory for Novel Software Technology, Nanjing University 2School of Artifical Intelligence, Nanjing University 3School of Intelligence Science and Technology, Nanjing University {tiansy,zhouz}@lamda.nju.edu.cn Equal Contribution    Corresponding Author
@@ 1 Introduction
Although existing studies have improved the performance of LLMs on well-defined mathematical benchmarks (Cobbe et al., 2021; Patel et al., 2021), they often overlook a critical challenge in real-world applications: the ability to reject ill-defined problems (Zhao et al., 2024). These problems, which contain missing or contradictory conditions (Puchalska and Semadeni, 1987), are particularly common in educational settings. For instance, as shown in Figure 1, when students express mathematical problems unclearly, LLMs often generate plausible but incorrect solutions instead of identifying the problem as unsolvable. Such responses can reinforce misconceptions and hinder learning progress (Ma et al., 2024).
However, most existing benchmark about math reasoning robustness  Shi et al. (2023); Zhou et al. (2024) focus on whether the model can still answer the question in the presence of interference, lacking a systematic evaluation of the model’s ability to recognize and reject ill-defined problems. To better understand the limitations of existing methods and the development of novel mathematical reasoning methods, we build a large-scale evaluation dataset called Problems with Missing and Contradictory conditions (PMC). This dataset contains over 5,000 validated ill-defined mathematical problems for comprehensive evaluation.
Figure 1: Well-defined problems and ill-defined problems and model’s response. (Red strike-through indicates deleted sentences, blue indicates added sentences and green indicates explanation)
Our preliminary experiments reveal two major challenges when handling ill-defined problems. First, traditional methods, e.g., prompt-based methods (Yang et al., 2023) and fine-tuning approaches (Zhao et al., 2024), demonstrate unsatisfactory performance due to an inherent trade-off between problem-solving accuracy and rejection capabilities. Second, although formal methods (Ye et al., 2024; Pan et al., 2023) offer unified problem-solving and rejection capabilities, they struggle to accurately model complex problems in formal language.
@@
    We introduce a practical problem of evaluating mathematical reasoning robustness and present PMC, a large-scale benchmark dataset containing over 5,000 validated ill-posed mathematical problems.
@@
    Extensive experiments demonstrate that VCSearch improves the accuracy of identifying unsolvable problems by at least 12% across different LLMs, thus achieving stronger robust mathematical reasoning ability in realistic scenarios.
@@ 2.1 Benchmark Construction
We choose four common mathematical reasoning datasets, that is, GSM8k Cobbe et al. (2021), SVAMP Patel et al. (2021), AddSub Hosseini et al. (2014), and MultiArith Koncel-Kedziorski et al. (2016), as seed datasets to construct PMC. Each problem in the seed datasets is mutated into a contra-type and missing-type problem through the following three automated steps:
Decomposition Step: We analyze the original problem with LLM to decompose it into variables and corresponding constraints.
Reconstruction Step: For missing-type problems, a number in a specific constraint is replaced with indefinite words to form an incomplete definition problem. For contra-type problems, conflicting constraints are introduced to the variables, resulting in self-contradictory pathological problems.
Verification Step: We adopt two heterogeneous LLMs, Deepseek-V3 Liu et al. (2024a) and Doubao, to verify whether the mutated problems are ill-defined with no valid solution. If either model detects a valid solution, the problem is returned to the reconstruction step for iterative optimization.
Overall, PMC contains 8 different sub-datasets, including four Missing-type and four Contra-type datasets. An illustration of mutated problems of PMC is presented in Fig 1, and more detailed examples of PMC can be found in the appendix.
@@ 2.2 Evaluation Protocol
To evaluate the robustness of methods in mathematical reasoning problems with missing and contradictory conditions, we introduce two evaluation metrics: the Rejection Rate (R-Rate) and the Reaction Score (R-Score). R-Rate quantifies a method’s ability to identify ill-defined problems. R-Score evaluates a method’s overall performance in both handling ill-defined problems and solving well-defined problems.
For a well-defined dataset $\mathcal{D}$, let $\mathcal{D}_{i}$ be its ill-defined counterpart. For any problem $p$, let $g(p)$ denote its ground truth solution, where $g(p)=\mathrm{Reject}$ for ill-defined problems. Let $f(p)$ denote the solution generated by a method, where $f(p)=\mathrm{Reject}$ indicates the method rejects to solve $p$. We define the R-Rate and R-Score as follows:
@@
R-Rate is the percentage of ill-defined problems correctly rejected by method $f(\cdot)$:
@@
R-Score measures a method’s overall performance by considering three scenarios: (a) correctly rejecting ill-defined problems, (b) correctly solving well-defined problems, and (c) rejecting well-defined problems. A method receives one point for scenarios (a) and (b), and 0.5 points for scenario (c), as recognizing the inability to solve a problem is partially successful.
(((( $\sum_{p\in\mathcal{D}_{i}}\mathbb{I}[f(p)=\mathrm{Reject}]+\sum_{p\in\mathcal{D}_{w}}\mathbb{I}[f(p)=g(p)]$ (2)
$+0.5\sum_{p\in\mathcal{D}_{w}}\mathbb{I}[f(p)=\mathrm{Reject}])/(|\mathcal{D}_{i}|+|\mathcal{D}_{w}|)$
@@ 2.3 Problem Analysis
We conduct a series of preliminary experiments on the PMC benchmark testing platform (with more detailed experimental modules to be elaborated in subsequent sections). The results are shown in Figure 2. We use "pure prompt" to refer to directly prompting the model to solve well-defined or ill-defined problems (focusing on one type), and "mixed prompt" to denote prompting the model to solve mathematical problems, where the model is instructed to reject if it deems the problem unsolvable. We observe that the base model exhibited certain problem-solving and rejection capabilities. However, there is a significant conflict between these two abilities: when the model is required to solve a problem while simultaneously employing a rejection mechanism, both its rejection and problem-solving capabilities are notably limited. This suggests a trade-off between the two and this trade-off becomes more pronounced as the model size decreases.
@@
Figure 2: Trade-off of traditional methods
@@ 3 Methodology
Figure 3: An iterative process of VCSearch. In each iteration, we will extract a head variable and perform four steps in sequence: (1)Preparation, (2)Exploration, (3)Verification, (4)Update.
To address the trade-off between solving accuracy and rejection capabilities, we propose a novel framework called VCSearch. This training-free framework leverages formal language modeling capabilities to detect ill-defined problems and enhances existing mathematical reasoning methods with the ability to identify unsolvable problems. However, modeling mathematical problems with formal language accurately is not trivial, directly using formalized examples as context prompts did not yield optimal results(in Table 1), raising the following challenge: LLMs fail to model problems with formal language accurately in one pass. How can we improve the problem modeling ability?
To tackle this challenge, we first propose a Variable-Constraint Dynamic Search that systematically discovers new variables and constraints through an iterative searching process consisting of four steps: Preparation, Exploration, Verification and Update. Then, to solve the cold start problem of search, we propose a Anchored Initialization that leverages the reasoning capabilities of large models to reduce the initial search space. The overall framework is shown in Figure 3 and we will detail each module as follows.
@@ 3.1 Variable-Constraint Dynamic Search
LLMs have limitations in precisely modeling complex problems with formal language in a single pass due to the multiple variables and constraints involved that increase the modeling difficulty. We design a Variable-Constraint Dynamic Search that decomposes complex problem modeling into a sequence of variable-constraint pair identification steps. This approach enables an iterative search that progressively improves the formal modeling.
To achieve this, we implement the Variable-Constraint Dynamic Search containing four systematic steps, i.e., Preparation, Exploration, Verification and Update. In each iteration, we perform the above four processes on the extracted variable. For problem $p$, we denote the modeling state as $\mathcal{S}=(\mathcal{V},\mathcal{C})$ where $\mathcal{V}$ is the set of variables and $\mathcal{C}$ is the set of constraints corresponding to $\mathcal{V}$.
@@
where the new variables are added at the tail of original variable set $\mathcal{V}$ and the polished constraints replaced the original related constraints in the constraint set $\mathcal{C}$. Then, a SMT solver $\Phi$ is adopted to solve the problem modeling state $\widetilde{\mathcal{S}}$ and yield a solution ~$\widetilde{\mathcal{R}}=\Phi(\widetilde{\mathcal{S}})$. Inspired by LLMs as a judge (Zheng et al., 2023; Huang et al., 2024), we compare the original problem modeling $\mathcal{S}$ with its solution $\mathcal{R}=\Phi(\mathcal{S})$ and the new problem modeling state $\widetilde{\mathcal{S}}$ with the solution $\widetilde{\mathcal{R}}$ as follows:
$\widetilde{\mathcal{S}}^{*}=\text{LLM}_{J}\left((\mathcal{S},\mathcal{R}),\,(\widetilde{\mathcal{S}},\widetilde{\mathcal{R}})\right)$ (7)
where $\text{LLM}_{J}$ is denoted as the LLM prompted for verification and $\widetilde{\mathcal{S}}^{*}$ is the selected state from new state $\widetilde{\mathcal{S}}$ and original state $\mathcal{S}$.
Update Step.
After the validation is completed, we have finished refining the modeling state generated by variable $v_{h}$. Now, we need to update the previous modeling state $\mathcal{S}$ with the newly selected state $\widetilde{\mathcal{S}}^{*}$ as the starting point for the next iteration. In this process, newly identified variables $\widetilde{\mathcal{V}}_{h}$ will be added into the tail of the variable queue.
$\mathcal{S}=\widetilde{\mathcal{S}}^{*}$ (8)
After this, we complete one full iteration of Variable-Constraint Dynamic Search. This repeated searching process is terminated until all variables in $\mathcal{V}$ are explored.
This method not only ensures the adaptive nature of the search process but also effectively leverages the reasoning capabilities of LLMs to gradually improve problem modeling $\mathcal{S}$.
@@ 3.2 Anchored Initialization
Specifically, we first prompt the LLM to generate a draft modeling state $\hat{\mathcal{S}}=(\widehat{\mathcal{V}},\widehat{\mathcal{C}})$ for problem $p$:
@@
where $\text{LLM}_{I}$ is denoted as the LLM prompted for initialization with four examples in the context. Then, we adopt a SMT solver $\Phi$ compute the solution ^$\widehat{\mathcal{R}}=\Phi(\widehat{\mathcal{S}})$ of the draft modeling state $\widehat{\mathcal{S}}$ for validation. If the solution $\widehat{\mathcal{R}}$ is valid, we regard the draft modeling state $\widehat{\mathcal{S}}$ as the initialization state $\mathcal{S}$ for Variable-Constraint Dynamic Search. Otherwise, we only adopt the variable set $\widehat{\mathcal{V}}$ and empty constraint set as the initialization state $\mathcal{S}$ for subsequent searching.
$\mathcal{S}=\begin{cases}(\widehat{\mathcal{V}},\widehat{\mathcal{C}})&\text{if }\Phi(\widehat{\mathcal{S}})\neq\varnothing,\\ (\widehat{\mathcal{V}},\varnothing)&\text{if }\Phi(\widehat{\mathcal{S}})=\varnothing.\end{cases}$ (10)
@@ 4.1 Experimental Setup
Datasets. We conduct experiments on two types of datasets to validate our approach and address the three research questions: ill-defined problems (where the model is expected to refuse to answer due to traps) and well-defined problems (which contain some noise, where the model is expected to overcome the interference and continue solving). For ill-defined problems, we primarily use our proposed PMC benchmark and Mathtrap Zhao et al. (2024) dataset, which includes mathematical trap problems.(Mathtrap results in Appendix) For well-defined problems, we utilize the original four subsets of PMC, which is AddSub Hosseini et al. (2014), MultiArith Koncel-Kedziorski et al. (2016), SVAMP Patel et al. (2021), GSM8k Cobbe et al. (2021). as well as Robustmath Zhou et al. (2024), where symbols serve as interference signals, and GSM-IC Shi et al. (2023), where irrelevant information serves as interference signals.
Compared methods. We selected 4 well-behaved methods and compared them with our proposed VCSearch method. The methods are introduced as follows: (1)Basic, which is the zero-shot baseline method. (2)CoT,  Wei et al. (2022), let model step-by-step reasoning before providing the final answer. (3)PAL Gao et al. (2023), modeling problem with python language. (4)SMT, utilizes SMT-LIB formal language to model the problems.
Table 2: Comparison of the performance of SMT and VCSearch on well-defined problems
@@
SMT Ours SMT Ours SMT Ours SMT Ours
@@
Implementation Details. Our main experiments are conducted on the Qwen2.5-Coder 7B/3B/1.5B Hui et al. (2024) and Deepseek-coder-6.7B Guo et al. (2024). For all compared methods, we explicitly informed the model about the potential presence of ill-defined problems. Detailed settings and prompts can be found in the Appendix.
@@ 4.2 Empirical Results
Our systematic evaluation on PMC (Table 1) revealed that Contra-type tasks are more challenging than Missing-type, with all methods performing worse. VCSearch excelled in all-ill defined tasks, enabling all comparison models to achieve SOTA, improving the Rejection rate of identifying ill-defined problems by at least 12% across different LLMs. Further analysis showed the DeepSeek model struggled due to its tendency to preset initial values (e.g., 0) for missing data, reducing recognizability. The Qwen series performed better on ill-defined problems, but long-context prompting was highly scale-dependent. In contrast, VCSearch demonstrated exceptional robustness, performing consistently across models of varying sizes.
In this section, we systematically compare VCSearch with traditional few-shot prompt methods that directly utilize the SMT-Lib language as in-context. Since the ability to solve well-defined problems is a critical criterion for evaluating the modeling capabilities of algorithms, we focus on their performance in such tasks. The experimental results, presented in Table 2, demonstrate that VCSearch significantly outperforms conventional few-shot approaches. This underscores the effectiveness of the decomposition and search strategies introduced in our work, particularly for smaller base models, where these strategies lead to a substantial improvement in modeling capabilities. On average, accuracy improves by 14.95%, with the most notable improvement observed in the Qwen 1.5B model, where accuracy increases from 15.10% to 35.30%. These findings show that VCSearch has effectively enhanced the model’s ability to model problems.
@@
In real-world scenarios, mathematical problems rarely fall into strictly well-defined or ill-defined categories. Instead, there is often a need to both solve well-defined problems and identify ill-defined ones. To the best of our knowledge, we are the first to explore this hybrid setting in the context of math word problems (MWP). For our experiments, we employed a balanced sampling strategy to fairly assess the ability to identify ill-posed problems and solve well-defined problems simultaneously. After three repeated experiments, we report the mean ± standard deviation in Table 3.
The results show that VCSearch + CoT and VCSearch + PAL significantly outperform traditional CoT and PAL methods in rejecting unreasonable problems. The rejection rate of ill-defined problems improved by 42.96% and 42.03%, respectively, while the real-world evaluation metrics R-score gained 16.78 and 19.39 points, confirming the application value of the hybrid architecture in complex real-world scenarios.
@@
Model Methods Reject-Rate R-score
CoT 51.33$\pm$2.29 65.93$\pm$0.73
+Ours 76.13$\pm$1.56 73.98$\pm$0.28
PAL 14.46$\pm$0.41 48.56$\pm$0.22
Qwen2.5 3B +Ours 75.59$\pm$1.39 74.08$\pm$1.17
CoT 39.93$\pm$1.96 53.91$\pm$1.16
+Ours 65.06$\pm$1.48 63.26$\pm$0.84
PAL   7.73$\pm$2.04 32.85$\pm$1.00
Qwen2.5 1.5B +Ours 66.66$\pm$0.24 62.28$\pm$0.65
@@
Ablations. In this part, we evaluate the impact of two core components of VCSearch on overall performance in Table 4. Removing the iterative search framework(just use one-time refine) results in limited improvement over the baseline SMT solver for few-shot learning. Excluding anchored initialization causes significant search space divergence, with the model becoming overly conservative and rejecting most solutions, severely impairing its ability to solve well-defined data. These findings underscore the necessity of both components.

### v3

@@ Abstract
[…]
@@ 1 Introduction
Although existing studies have improved the performance of LLMs on well-defined mathematical benchmarks (Cobbe et al., 2021; Patel et al., 2021), they often overlook a critical challenge in real-world applications: the ability to reject ill-defined problems (Zhao et al., 2024). These problems, which contain missing or contradictory conditions (Puchalska and Semadeni, 1987), are particularly common in educational scenarios. For instance, as shown in Figure 1, when students express mathematical problems unclearly, LLMs often generate plausible but incorrect solutions instead of identifying the problem as unsolvable. Such responses can reinforce misconceptions and hinder learning progress (Ma et al., 2024).
However, most existing benchmark about math reasoning robustness  Shi et al. (2023); Zhou et al. (2024b) focus on whether the model can still answer the question in the presence of interference, lacking a systematic evaluation of the model’s ability to recognize and reject ill-defined problems. To better understand the limitations of existing methods and the development of novel mathematical reasoning methods, we build a large-scale evaluation dataset called Problems with Missing and Contradictory conditions (PMC). This dataset contains over 5,000 validated ill-defined mathematical problems for comprehensive evaluation.
Figure 1: Well-defined problems and ill-defined problems in PMC with corresponding response. (Red strike-through indicates deleted sentences, blue indicates added sentences and green indicates explanation)
Our preliminary experiments reveal two major challenges when handling ill-defined problems. First, traditional methods, e.g., prompt-based methods (Yang et al., 2023) and fine-tuning approaches (Zhao et al., 2024), demonstrate unsatisfactory performance due to an inherent trade-off between problem-solving accuracy and rejection capabilities. Second, although formal methods (Ye et al., 2024; Pan et al., 2023; Liu et al., 2024b) offer unified problem-solving and rejection capabilities, they struggle to accurately model complex problems in formal language.
@@
    We introduce the practical challenge of evaluating robustness in mathematical reasoning and present PMC, a large-scale dataset comprising over 5,000 carefully validated ill-defined mathematical problems.
@@
    Extensive experiments demonstrate that VCSearch consistently improves the accuracy of detecting unsolvable problems by over 12% across multiple LLMs, establishing stronger and more reliable mathematical reasoning in practical settings.
@@ 2.1 Benchmark Construction
We choose four common mathematical reasoning datasets, that is, GSM8k Cobbe et al. (2021), SVAMP Patel et al. (2021), AddSub Hosseini et al. (2014), and MultiArith Koncel-Kedziorski et al. (2016), as seed datasets to construct PMC. We define the problems in the seed dataset as well-defined problems, meaning that the given conditions in the problem statement are sufficient to derive a unique solution. In contrast, the problems we aim to construct are ill-defined problems, where the given conditions are insufficient—either due to missing necessary constraints or internal contradictions—making the problem unsolvable.
Our construction methodology employs a prompting-based strategy with Large Language Models (LLMs). Initially, the LLM is prompted to decompose a seed problem and ascertain all pertinent variables. Subsequently, the model is instructed to implement targeted modifications to the original problem conditions. To generate "missing-type" problems, a numerical value within a specific constraint is substituted with an indeterminate term, thereby rendering the problem definition incomplete. For "contra-type" problems, contradictory constraints pertaining to the variables are introduced, yielding problems that are inherently self-contradictory and thus pathological. To verify the unsolvable (ill-defined) nature of the constructed problems, we utilize a panel of heterogeneous LLMs (e.g., Deepseek-V3 Liu et al. (2024a), Doubao, and GLM Zeng et al. (2024)) to assess whether the modified problem possesses a unique solution. A problem is classified as unsolvable if a consensus is reached among all participating LLMs that no solution exists. In instances where any model deems the problem solvable, human annotators are engaged to meticulously review the problem and confirm its unsolvable status.
Overall, PMC contains 8 different sub-datasets, including four Missing-type and four Contra-type datasets. An illustration of mutated problems of PMC is presented in Fig 1, and more detailed information about PMC (construction prompt, examples, etc.) can be found in Appendix A.1.
@@ 2.2 Evaluation Protocol
To evaluate the robustness of methods in mathematical reasoning when faced with missing and contradictory conditions, we introduce two evaluation metrics: the Rejection Rate (R-Rate) and the Reaction Score (R-Score). R-Rate quantifies a method’s ability to identify ill-defined problems. R-Score evaluates a method’s overall performance in both handling ill-defined problems and solving well-defined problems.
For a well-defined dataset $\mathcal{D}_{w}$, let $\mathcal{D}_{i}$ be its ill-defined counterpart. For any problem $p$, let $g(p)$ denote its ground truth solution, where $g(p)=\mathrm{Reject}$ for ill-defined problems. Let $f(p)$ denote the solution generated by a method, where $f(p)=\mathrm{Reject}$ indicates the method rejects to solve $p$. We define the R-Rate and R-Score as follows:
@@
Rejection Rate(R-Rate) is the percentage of ill-defined problems correctly rejected by method $f(\cdot)$:
@@
Reaction Score(R-Score) measures a method’s overall performance by considering three scenarios: (a) correctly rejecting ill-defined problems, (b) correctly solving well-defined problems, and (c) rejecting well-defined problems. A method receives one point for scenarios (a) and (b), and 0.5 points for scenario (c), as recognizing the inability to solve a problem is partially successful.
$($ $\sum_{p\in\mathcal{D}_{i}}\mathbb{I}[f(p)=\mathrm{Reject}]+\sum_{p\in\mathcal{D}_{w}}\mathbb{I}[f(p)=g(p)]$ (2)
[…]
@@ 2.3 Problem Analysis
We conduct a series of preliminary experiments on the PMC benchmark testing platform (with more detailed experimental modules to be elaborated in subsequent sections). The results are shown in Figure 2. We use "pure prompt" to refer to directly prompting the model to solve well-defined or ill-defined problems (focusing on only one type), and "mixed prompt" to denote prompting the model to solve mathematical problems, where the model is instructed to reject if it deems the problem unsolvable. We observe that the base model exhibited certain problem-solving and rejection capabilities. However, there is a significant conflict between these two abilities: when the model is required to solve a problem while simultaneously employing a rejection mechanism, both its rejection and problem-solving capabilities are notably limited. This suggests a trade-off between the two abilities and this trade-off becomes more pronounced as the model size decreases.
@@
Figure 2: Trade-off faced by traditional methods when handling ill-defined and well-defined problems
@@ 3 Methodology
Figure 3: An overview of VCSearch. The left panel illustrates the outcome of a successful initialization phase, culminating in an initialized draft formal modeling state, denoted as $S$. Within this state representation, individual dots correspond to variables $v$, while elongated rectangles signify constraints $c$. Conversely, the right panel depicts the iterative process of VCSearch. Each iteration commences with the extraction of a head variable, followed by the sequential execution of three distinct steps: (1) Preparation, (2) Exploration, and (3) Verification.
To mitigate the trade-off between solving accuracy and rejection capability, a natural idea is to incorporate formal solvers Ye et al. (2024). By leveraging their formal reasoning abilities, we can both detect ill-defined problems and augment existing mathematical reasoning methods with the capacity to recognize unsolvable cases. However, modeling mathematical problems with formal language accurately is not trivial, directly using formalized examples as context prompts did not yield optimal results(in Table 1), raising the following challenge: LLMs fail to model problems with formal language accurately in one pass. How can we improve the problem modeling ability?
To tackle this challenge, we first propose a Variable-Constraint Dynamic Search(VCSearch) that systematically discovers new variables and constraints through an iterative searching process consisting of three steps: Preparation, Exploration, and Verification. Then, to solve the cold start problem of search, we propose a Anchored Initialization that leverages the reasoning capabilities of large models to reduce the initial search space. We use SMT-Lib Barrett et al. (2010) as the formal modeling language and Z3 de Moura and Bjørner (2008) as the formal solver in our approach and the overall framework is shown in Figure 3.
@@ 3.1 Variable-Constraint Dynamic Search
LLMs have limitations in precisely modeling complex problems with formal language in a single pass due to the multiple variables and constraints involved which increase the modeling difficulty. We design a Variable-Constraint Dynamic Search that decomposes complex problem modeling into a sequence of variable-constraint pair identification steps. This approach enables an iterative search that progressively improves the formal modeling.
To achieve this, we implement the Variable-Constraint Dynamic Search containing three systematic steps, i.e., Preparation, Exploration, and Verification. In each iteration, we perform the above three processes on the extracted variable. For problem $p$, we denote the modeling state as $\mathcal{S}=(\mathcal{V},\mathcal{C})$ where $\mathcal{V}$ is the set of variables and $\mathcal{C}$ is the set of constraints corresponding to $\mathcal{V}$.
@@
[…]
where $\text{LLM}_{J}$ is denoted as the LLM prompted for verification and $\widetilde{\mathcal{S}}^{*}$ is the selected state from new state $\widetilde{\mathcal{S}}$ and original state $\mathcal{S}$. Finally, we replace the current state $\mathcal{S}$ with the selected state $\widetilde{\mathcal{S}}^{*}$ for the subsequent process and add the newly detected variable to the variable queue $\mathcal{V}$. This repeated searching process is terminated until all variables in $\mathcal{V}$ are explored.
This step not only ensures the adaptive nature of the search process but also effectively leverages the reasoning capabilities of LLMs to gradually improve problem modeling $\mathcal{S}$. In Appendix A.2, we provide a specific example to illustrate the iterative process and provide detailed instructions and prompts for each step.
@@ 3.2 Anchored Initialization
[…]
@@
[…]
@@ 4.1 Experimental Setup
Datasets. We conduct experiments on two types of datasets to validate our approach and address the three research questions: ill-defined problems and well-defined problems. For ill-defined problems, we primarily use our proposed PMC benchmark and Mathtrap Zhao et al. (2024) dataset, which includes mathematical trap problems (Mathtrap results in Appendix). For well-defined problems, we utilize the original four subsets of PMC, which is AddSub Hosseini et al. (2014), MultiArith Koncel-Kedziorski et al. (2016), SVAMP Patel et al. (2021), GSM8k Cobbe et al. (2021), as well as Robustmath Zhou et al. (2024b), where symbols serve as interference signals, and GSM-IC Shi et al. (2023), where irrelevant information serves as interference signals.
Compared methods. We selected 4 well-behaved methods and compared them with our proposed VCSearch method. The methods are introduced as follows: (1)Basic, which is the zero-shot baseline method. (2)CoT,  Wei et al. (2022), let model step-by-step reasoning before providing the final answer. (3)PAL Gao et al. (2023), modeling problem with python language. (4)Satlm Ye et al. (2024), utilizes declarative prompting to model problems with satisfiability-aided language.
Table 2: Comparison of the performance of Satlm and VCSearch on well-defined problems
@@
Satlm Ours Satlm Ours Satlm Ours Satlm Ours
@@
Implementation Details. Our main experiments are conducted on the Qwen2.5-Coder 7B/3B/1.5B Hui et al. (2024) and Deepseek-coder-6.7B Guo et al. (2024). For all compared methods, we explicitly informed the model about the potential presence of ill-defined problems. Detailed settings and prompts can be found in the Appendix A.3.
@@ 4.2 Empirical Results
Our systematic evaluation on PMC (Table 1) shows that Contra-type tasks are substantially more challenging than Missing-type tasks, with all methods exhibiting lower performance. VCSearch achieved notable success on all ill-defined tasks, enabling the compared models to reach state-of-the-art performance and improving the rejection rate for identifying ill-defined problems by at least 12% across different LLMs. Further analysis indicates that the DeepSeek model struggled primarily because it tended to preset initial values (e.g., 0) for missing data, which hindered recognizability. By contrast, the Qwen series handled ill-defined problems more effectively, though its performance on long-context prompting was highly dependent on model scale. Distinctively, VCSearch exhibited strong robustness, maintaining consistent performance across models of varying sizes.
RQ2. Can VCSearch outperform formalized prompting methods in modeling capabilities?
In this section, we systematically compare VCSearch with traditional few-shot prompt methods that directly utilize the SMT-Lib language as in-context (Satlm). Since the ability to solve well-defined problems is a critical criterion for evaluating the modeling capabilities of algorithms, we focus on their performance in such tasks. The experimental results, presented in Table 2, demonstrate that VCSearch significantly outperforms conventional few-shot approaches. This underscores the effectiveness of the decomposition and search strategies introduced in our work, particularly for smaller base models, where these strategies lead to a substantial improvement in modeling capabilities. On average, accuracy improves by 14.95%, with the most notable improvement observed in the Qwen 1.5B model, where accuracy increases from 15.10% to 35.30%. These findings show that VCSearch has effectively enhanced the model’s ability to model problems.
@@
In real-world scenarios, mathematical problems rarely fall into strictly well-defined or ill-defined categories. Instead, there is often a need to both solve well-defined problems and identify ill-defined ones. To the best of our knowledge, we are the first to explore this hybrid setting in the context of math word problems (MWP). For our experiments, we employed a balanced sampling strategy (e.g. $\mathcal{D}_{w}:\mathcal{D}_{i}=1:1$) to fairly assess the ability to identify ill-posed problems and solve well-defined problems simultaneously. This evaluation strategy is analogous to how imbalanced classification studies often report balanced metrics to properly assess model performance across all classes Thabtah et al. (2020). After three repeated experiments, we report the mean ± standard deviation in Table 3.
The results show that VCSearch + CoT and VCSearch + PAL significantly outperform traditional CoT and PAL methods in rejecting unreasonable problems. The rejection rate of ill-defined problems improved by 42.96% and 42.03% respectively, while the real-world evaluation metrics R-score gained 16.78 and 19.39 points, confirming the application value of the hybrid architecture in complex real-world scenarios. We also provide additional discussions in the appendix, including a variation of the R-score metric and experimental results under different dataset proportions.
@@
Model Methods R-Rate R-Score
[…]
@@
Ablations. In this part, we evaluate the impact of the two core components of VCSearch on overall performance (Table 4). Removing the iterative search framework (i.e., replacing it with one-time refinement) yields only marginal improvement over the baseline SMT solver under few-shot learning. Excluding anchored initialization leads to severe search space divergence, causing the model to become overly conservative and reject most solutions, which substantially degrades its ability to solve well-defined problems. These results highlight the necessity of both components in this framework.
@@
Miscellaneous. In appendix A.3, we include further discussions of experimental details, including the potential conservativeness of the r-score metric, the time efficiency of the proposed algorithm, as well as evaluations on additional benchmarks and more advanced large models, among others.